Circular No. 6313

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

R CORONAE BOREALIS Y. Efimov, Crimean Astrophysical Observatory (CrAO); and A. E.

Rosenbush, Main Astronomical Observatory, Kiev, report data on the photopolarimetry of R CrB, which was observed at deep minimum. The UBVRI observations were made with 1.25-m CrAO reflector, given here as magnitude (± 0.01), percentage polarization, and position angle in degrees: Jan. 2.1073 UT, U = 13.7, 4.85 ± 0.37, 98^{o}.0 ± 2^{o}.2; B = 14.09, 6.02 ± 0.32, 105^{o}.8 ± 1^{o}.5; V = 13.61, 5.45 ± 0.52, 108^{o}.8 ± 2^{o}.7; R = 13.07, 4.59 ± 0.21, 116^{o}.0 ± 1^{o}.3; I = 12.10, 2.19 ± 0.24, 112^{o}.7 ± 3^{o}.1. Jan. 26.0802, U = –, 3.50 ± 0.49, 107^{o}.1 ± 4^{o}.0; B = –, 6.03 ± 0.31, 106^{o}.3 ± 1^{o}.5; V = –, 3.47 ± 0.73, 105^{o}.5 ± 6^{o}.0; R = –, 4.00 ± 0.28, 121^{o}.0 ± 2^{o}.0; I = –, 2.18 ± 0.28, 126^{o}.8 ± 3^{o}.6. Jan. 27.0928, U = 13.55, 3.93 ± 0.29, 103^{o}.1 ± 2^{o}.1; B = 13.97, 4.97 ± 0.27, 105^{o}.3 ± 1^{o}.6; V = 13.50, 5.53 ± 0.43, 115^{o}.0 ± 2^{o}.2; R = 13.02, 4.22 ± 0.16, 120^{o}.1 ± 1^{o}.1; I = 11.95, 1.89 ± 0.15, 126^{o}.0 ± 2^{o}.3.

GRO J1744–28 Further to their items on IAUC 6309 and 6310, T. Augusteijn et al.

report: “Due to an error in the astrometric analysis the star reported on those Circulars is offset from the radio source in the XTE error box by 10′′ and is therefore unrelated to the radio source.”

3C 279 A. R. Cooray and A. B. Fletcher, Massachusetts Institute of Technol-

ogy, write: “The recent γ-ray-brightening quasar 3C 279 (IAUC 6294; z = 0.57) was observed during Feb. 11–14 with the Cerro Tololo 1.5-m telescope. The magnitudes during this 4-day period remained steadily at V = 15.5 and R = 15.0; no nightly variation was found at the accuracy of our present photometry (∼ 0.1 mag).”

COMET C/1995 Y1 (HYAKUTAKE) D. Schleicher, Lowell Observatory, reports: “Narrow-band photometry

of comet C/1995 Y1 was obtained with the Lowell Observatory 0.8-m telescope on Feb. 9.5 UT. Haser-model production rates were: log Q(OH) = 28.40, log Q(CN) = 25.76, log Q(C_{2}) = 25.81. The dust production rate was log (Afρ) = 2.34.”

1996 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green